# Parabolic bundles and the intersection cohomology of moduli spaces of vector bundles on curves

### v1

@@ 3.2. Local structure of $\mathcal{M}_{0}(r)$
(17) $\mathcal{M}_{0}(r)=\mathcal{M}_{0}(r)^{{\mathrm{ab}}}\bigsqcup\mathcal{M}_{0}(r)^{{\mathrm{nab}}}\quad\text{and }M_{\rho}=M_{\rho}^{\mathrm{ab}}\bigsqcup M_{\rho}^{\mathrm{nab}}.$
@@
Now we recall some well known results on the infinitesimal structure of the moduli space $\mathcal{M}_{0}(r)$. For the proofs, we refer the reader to [22, 28].
@@
Rescaling $\mathrm{Ext}^{1}(V,V)$ by $\lambda\in\mathbb{C}^{*}$ commutes with the action of $\mathrm{Aut}(V)$, and thus descends as an action of $\mathbb{C}^{*}$ on $U_{V}$.
Now let $V\cong V_{1}\oplus\ldots\oplus V_{k}\in M_{\rho}^{\mathrm{ab}}$ be an abelian point.; then
@@
  • $\mathrm{Aut}(V)$ is a torus $(\mathbb{C}^{*})^{k}$ with an element $(t_{1},\ldots,t_{k})\in\left(\mathbb{C}^{*}\right)^{k}$ acting diagonally on $V_{1}\oplus\ldots\oplus V_{k}$;
@@
Hence we can describe the local neighborhood $U_{V}$ of $V$ in $\mathcal{M}_{0}(r)$ as follows (see [31, §2]).
@@ 3.3. Moduli spaces of parabolic bundles
In this subsection, we recall some basic definitions and facts about parabolic bundles on curves. We will only present a simplified theory, necessary for this paper. For a complete description we refer to [24, 4, 32].
  • Fix a point $p\in{C}$. A parabolic bundle on $C$ is a vector bundle $W$ with a subspace $F_{1}\subset W_{p}$ in the fiber over $p$:
@@
If $W^{\prime}\subset{W}$ is a subbundle of $W$, then both $W^{\prime}$ and the quotient $W/W^{\prime}$ inherit a parabolic structure from $W$ in a natural way (cf. [24, Definition 1.7]).
@@
$c=(c_{1},\ldots,c_{1},c_{2},\ldots,c_{2})$
for the sequence with $r$ elements, where each weight $c_{i}$ is repeated $f_{i}$ times. Then without loss of generality, we can assume that the parabolic weights of a rank-$r$ and degree-$d$ bundle belong to the set
@@
We call a vector $c=(c_{1},\ldots,c_{r})\in\mathbb{R}^{r}$ such that $\sum_{i}c_{i}\in\mathbb{Z}$ regular if for any nontrivial subset $\Psi\subset\{1,2,\ldots,r\}$, we have $\sum_{i\in\Psi}c_{i}\notin\mathbb{Z}$.
For fixed rank $r$, degree $d$ and regular $c\in\Delta_{d}$, Mehta and Seshadri [24] constructed a smooth projective moduli space $\mathcal{P}_{d}(c)$ of stable parabolic bundles whose points are in one-to-one correspondence with the set of isomorphism classes of stable parabolic bundles of weight $c$.
@@
For a regular $c\in\Delta_{d}$, there exist universal bundles $U$ over $\mathcal{P}_{d}(c)\times C$ endowed with a flag $\mathcal{F}_{1}\subset\mathcal{F}_{2}=U_{p}$, and satisfying the standard tautological properties. In general, such universal bundles $U$ are unique only up to tensoring by the pull-back of a line bundle from $\mathcal{P}_{d}(c)$.
@@
We will say that a universal bundle $U$ over $\mathcal{P}_{0}(r)\times C$ is normalized if the line subbundle $\mathcal{F}_{1}\subset{U}_{p}$ is trivial.
@@
(ii) Let $(W,F_{1})$ be a stable parabolic bundle which represents a point in $\mathcal{P}_{1}(r)$. Then the vector bundle $W$ is stable, and thus represents a point in $\mathcal{M}_{1}(r)$. Moreover, $\mathcal{P}_{1}(r)$ is isomorphic to the projectivization of a universal bundle $U_{1}(r)_{p}$ over $\mathcal{M}_{1}(r)\times\{p\}$:
$\mathcal{P}_{1}(r)\cong\mathbb{P}(U_{1}(r)_{p})\xrightarrow{\pi_{1}}\mathcal{M}_{1}(r).$
@@ 3.4. The tautological Hecke correspondence
One can define (cf. [32, §7]) the tautological Hecke operators between the moduli spaces of parabolic bundles with different degrees and parabolic weights as follows. Given a vector bundle $W$ on $C$ with a flag $0=F_{0}\subsetneq F_{1}\subsetneq F_{2}=W_{p}$ in the fiber over $p\in C$, we consider the associated sheaf of sections $\mathcal{W}$ and define the subsheaf
@@
[32, Proposition 7.1] The map $\mathcal{H}$ induces an isomorphism of the moduli spaces
@@
3.5. Interlude on mixed Hodge structures and the Decomposition Theorem
For most of our arguments, the ”elementary” version of the Decomposition Theorem, Theorem 1.7 suffices, but there is an important technical point, which requires a refinement of this statement: we will need to know that the fibers of the map $\pi$ defined in (21) have pure Hodge structure. Below, we give the necessary definitions and background, but we will be brief, as will we only need a rather trivial part of the theory. We refer the reader to [9] and [8, 5] for more details.
3.5.1. Mixed Hodge structures
The $k$th cohomology of a smooth projective variety $Y$ carries a (pure) Hodge structure of weight $k$:
The cohomology of a not necessarily projective and not necessarily smooth varieties carry mixed Hodge structure (MHS): given such a variety $Y$, this is a pair $(F^{\bullet},W^{\bullet})$ where
When $Y$ is projective, we have $W_{k}\supset H^{k}(Y,\mathbb{C})$, and we say that the cohomology of $Y$ is pure if $W_{k-1}(H^{k}(Y,\mathbb{C}))=0$. A key, defining criterion of purity is the following:
Proposition 3.13.
A projective algebraic variety $Y$ has pure cohomology if and only if for a (and thus for any) birational map $\widetilde{Y}\to Y$ from a smooth projective variety $\widetilde{Y}$ to $Y$, the pullback map $H^{\bullet}(Y)\to H^{\bullet}(\widetilde{Y})$ is injective.
Some extensions of this notion:
  • Compactly supported cohomology $H^{\bullet}_{c}(Y)$ has MHS compatible with the MHS on cohomology.
  • MHS may be defined on intersection cohomologies and compactly supported intersection cohomologies as well.
  • Finally, one can extend the notion of MHS to intersection cohomologies with twisted coefficients (see §1.2), replacing the constant sheaf $\mathbb{Q}_{Y}$ with any local system $\mathcal{L}$ supported on an open dense subset of $Y$.
3.5.2. MHS and the Decomposition Theorem
It is a deep statement of Saito’s theory [8, 5, 29] that it is always possible to choose an isomorphism in the Decomposition Theorem 1.7 which respects the Mixed Hodge structures in (1.7.C) and (1.7.B). In particular, in (1.7.B), the MHS on the cohomology of the fiber $f^{-1}(p)$ is the direct sum of the MHS on the intersection cohomologies with twisted coefficients on the right hand side:
(22) $\mathcal{H}^{\bullet}\left(IC(\overline{Y}_{\alpha},\mathcal{L}_{\alpha})\right)_{p}\cong IH^{\bullet}(U_{\alpha},\mathcal{L}_{\alpha}),\quad U_{\alpha}\text{ is a neighborhood of }p\text{ in }\overline{Y}_{\alpha}.$
We will only need a small, rather trivial part of this theory, which will permit us to conclude that $H^{\bullet}(f^{-1}(p))$ has pure Hodge structure.
The exact statement is as follows:
Proposition 3.14.
Let $f:X\rightarrow Y$ be a proper algebraic map of algebraic varieties as in Theorem 1.7, and let $p\in Y$. Assume that for each stratum $Y_{\alpha}$ containing $p$ in its closure, the intersection cohomology $IH^{\bullet}(U_{\alpha})$ (see (22)) has pure Hodge structure, and that $\mathcal{L}_{\alpha}$ has trivial monodromy. Then the MHS on the cohomology of $H^{\bullet}(f^{-1}(p))$ is pure. More generally, the same conclusion holds if the monodromy of $\mathcal{L}_{\alpha}$ is finite, and the variation of Hodge structure defined by $\mathcal{L}_{\alpha}$ is trivial.
Next, we describe a simple situation, when the intersection cohomology of $IH^{\bullet}(U_{\alpha})$ is pure (see [23, Proposition 2.4]).
Definition 3.15.
  • the fixed point set is proper;
  • For every $p\in X$, the limit $\lim_{t\to 0}t\cdot p$ exists.
Proposition 3.16.
Assume that in the setup above, $U_{\alpha}$ is a semiprojective variety. Then the intersection cohomology $IH^{\bullet}(U_{\alpha})$ of $U_{\alpha}$ has pure Hodge structure.
Now, we observe that the $\mathbb{C}^{*}$ scaling action in Remark 3.6 contracts the étale neighbourhood $U_{V}$ to the origin, and this shows that $U_{V}$ is a semiprojective variety.
Corollary 3.17.
The intersection cohomology $IH^{\bullet}(U_{V})$ of $U_{V}$ introduced in Proposition 3.4 has pure Hodge structure.
3.5.3. Weight polynomials
Definition 3.18.
Let $Gr^{W}_{k}(H_{c}^{j}(Y)):=W_{k}(H_{c}^{j}(Y))/W_{k-1}(H_{c}^{j}(Y))\otimes\mathbb{C}$ be the weight-$k$ part of the $j$th compactly supported cohomology of $Y$. The weight polynomial of $Y$ is the generating function
$E_{t}(Y)=\sum_{j,k}(-1)^{j}\dim Gr^{W}_{k}(H_{c}^{j}(Y))t^{k}.$
Proposition 3.19.
If $Z\subset Y$ is a locally closed subvariety of the variety $Y$, then
$E_{t}(Y)=E_{t}(Z)+E_{t}(Y\setminus Z).$
@@ 3.6. Our plan
At this point, in order to write down (1.7.B) for our case, and to streamline our notation for the rest of the paper, we introduce the following conventions:
@@
At this point, we will be ready for the final step iii.; the equality (1.7.C) has the following form in our case:
$H^{\bullet}(\mathcal{P}_{0}(r))\cong\bigoplus_{\rho\vdash r}IH^{\bullet}(\overline{M}_{\rho},\mathcal{L}_{\rho}).$
Proposition 3.10 and Corollary 3.12 tell us that (as graded vector spaces). Finally, a recursion leading to our main result, Theorem 1.4 is developed in §8.

### v2

@@ 3.2. Local structure of $\mathcal{M}_{0}(r)$
(17) $\mathcal{M}_{0}(r)=\mathcal{M}_{0}(r)^{{\mathrm{ab}}}\sqcup\mathcal{M}_{0}(r)^{{\mathrm{nab}}}\quad\text{and }M_{\rho}=M_{\rho}^{\mathrm{ab}}\sqcup M_{\rho}^{\mathrm{nab}}.$
@@
Remark 3.4.
The locus $M_{\rho}^{\mathrm{ab}}$ associated with a partition $\rho=[r_{1},\ldots,r_{k}]\vdash r$ is the quotient of the variety
$\mathcal{M}_{\rho}^{\mathrm{ab}}=\{(V_{1},V_{2},\dots,V_{k})|\,V_{i}\in\mathcal{M}_{0}(r_{i})^{\mathrm{st}},V_{i}\not\cong V_{j}\text{ if }i\neq j\}\subset\\ \mathcal{M}_{0}(r_{1})^{\mathrm{st}}\times\mathcal{M}_{0}(r_{2})^{\mathrm{st}}\times\dots\times\mathcal{M}_{0}(r_{k})^{\mathrm{st}}$
by the action of the finite group $\mathrm{Aut}(\rho)\subset\Sigma_{k}$:
This cover extends to a finite morphism $g_{\rho}:\mathcal{M}_{0}(r_{1})\times\ldots\times\mathcal{M}_{0}(r_{k})\rightarrow\overline{M}_{\rho}$.
Now we recall some well-known results about the infinitesimal structure of the moduli space $\mathcal{M}_{0}(r)$. For the proofs, we refer the reader to [24, 30].
@@
On the other hand, the rescaling of $\mathrm{Ext}^{1}(V,V)$ by $\lambda\in\mathbb{C}^{*}$ commutes with the action of $\mathrm{Aut}(V)$, and thus descends to an action of $\mathbb{C}^{*}$ on $U_{V}$.
Definition 3.8.
  • the fixed point set is proper;
  • For every $p\in X$, the limit $\lim_{t\to 0}t\cdot p$ converges.
Clearly, we have:
Corollary 3.9.
Let $V\in\mathcal{M}_{0}(r)$. Then, endowed with the $\mathbb{C}^{*}$-action described in Remark 3.7, the local model $U_{V}$ is a semiprojective variety.
We will spell this out more explicitly in the case when $V\cong V_{1}\oplus\ldots\oplus V_{k}\in M_{\rho}^{\mathrm{ab}}$ is an abelian point. Then
@@
  • $\mathrm{Aut}(V)$ is a torus $(\mathbb{C}^{*})^{k}$ with $(t_{1},\ldots,t_{k})\in\left(\mathbb{C}^{*}\right)^{k}$ acting diagonally on $V_{1}\oplus\ldots\oplus V_{k}$;
@@
Hence we can describe the local neighborhood $U_{V}$ of $V$ in $\mathcal{M}_{0}(r)$ as follows (see [36, §2]).
@@ 3.3. Moduli spaces of parabolic bundles
In this paragraph, we recall some basic definitions and facts about parabolic bundles on curves. We will only present a simplified theory, necessary for this paper. For a complete description we refer to [26, 4, 35].
  • Fix a point $p\in{C}$. We define a parabolic bundle on $C$ as a vector bundle $W$ with a subspace $F_{1}\subset W_{p}$ in the fiber over $p$:
@@
If $W^{\prime}\subset{W}$ is a subbundle of $W$, then both $W^{\prime}$ and the quotient $W/W^{\prime}$ inherit a parabolic structure from $W$ in a natural way (cf. [26, Definition 1.7]).
@@
$\boldsymbol{c}=(c_{1},\ldots,c_{1},c_{2},\ldots,c_{2})$
for the sequence with $r$ elements, where each weight $c_{i}$ is repeated $f_{i}$ times. Then, without loss of generality, we can assume that the parabolic weights of a rank-$r$ and degree-$d$ bundle belong to the set
@@
We call a vector $\boldsymbol{c}=(\gamma_{1},\ldots,\gamma_{r})\in\mathbb{R}^{r}$ such that $\sum_{i}\gamma_{i}\in\mathbb{Z}$ regular if for any nontrivial subset $\Psi\subset\{1,2,\ldots,r\}$, we have $\sum_{i\in\Psi}\gamma_{i}\notin\mathbb{Z}$.
For fixed rank $r$, degree $d$ and regular $\boldsymbol{c}\in\Delta_{d}$, Mehta and Seshadri [26] constructed a smooth projective moduli space $\mathcal{P}_{d}(\boldsymbol{c})$ of stable parabolic bundles whose points are in one-to-one correspondence with the set of isomorphism classes of stable parabolic bundles of weight $\boldsymbol{c}$.
@@
For a regular $\boldsymbol{c}\in\Delta_{d}$, there exist universal bundles $\mathcal{U}$ over $\mathcal{P}_{d}(\boldsymbol{c})\times C$ endowed with a flag $\mathcal{F}_{1}\subset\mathcal{F}_{2}=\mathcal{U}_{p}$, and satisfying the standard tautological properties. In general, such universal bundles $\mathcal{U}$ are unique only up to tensoring by the pull-back of a line bundle from $\mathcal{P}_{d}(\boldsymbol{c})$.
@@
We will say that a universal bundle $\mathcal{U}$ over $\mathcal{P}_{0}(r)\times C$ is normalized if the line subbundle $\mathcal{F}_{1}\subset{\mathcal{U}}_{p}$ is trivial.
@@
(ii) Let $(W,F_{1})$ be a stable parabolic bundle which represents a point in $\mathcal{P}_{1}(r)$. Then the vector bundle $W$ is stable, and thus represents a point in $\mathcal{M}_{1}(r)$. Moreover, $\mathcal{P}_{1}(r)$ is isomorphic to the projectivization of a universal bundle $\mathcal{U}_{1}(r)_{p}$ over $\mathcal{M}_{1}(r)\times\{p\}$:
$\mathcal{P}_{1}(r)\cong\mathbb{P}(\mathcal{U}_{1}(r)_{p})\xrightarrow{\pi_{1}}\mathcal{M}_{1}(r).$
@@ 3.4. The tautological Hecke correspondence
One can define (cf. [35, §7]) the tautological Hecke operators between the moduli spaces of parabolic bundles with different degrees and parabolic weights as follows. Given a vector bundle $W$ on $C$ with a flag $0=F_{0}\subsetneq F_{1}\subsetneq F_{2}=W_{p}$ in the fiber over $p\in C$, we consider the associated sheaf of sections $\mathcal{W}$ and define the subsheaf
@@
[35, Proposition 7.1] The map $\mathcal{H}$ induces an isomorphism of the moduli spaces
@@ 3.5. Our plan
Motivated by this vanishing, in order to write down (1.7.B) for our case, and to streamline our notation for the rest of the paper, we introduce the following conventions:
@@
At this point, we will be ready for the final step iii.: the equality (1.7.C) has the following form in our case:
$H^{\bullet}(\mathcal{P}_{0}(r))\cong\bigoplus_{\rho\vdash r}IH^{\bullet}(\overline{M}_{\rho},\mathcal{L}_{\rho}),$
and, in §9, we develop this into a recursion leading to our main result, Theorem 1.4.
However, before we can embark on this road, we need to make small technical detour.
4. Interlude on mixed Hodge structures and the Decomposition Theorem
For most of our arguments, the ”elementary” version of the Decomposition Theorem, Theorem 1.7 suffices, but there is an important technical point, which requires a refinement of this statement: we will need to know that the fibers of the map $\pi$ defined in (21) have pure Hodge structure, and for this, we will need Saito’s version of the Decomposition Theorem, and a basic observation from the theory of Mixed Hodge modules [32].
Below, we give the necessary definitions and background, but we will be brief, as we will only need a rather trivial part of the theory. We refer the reader to [10] and [9, 6] for more details.
4.1. Mixed Hodge structures
The $k$-th cohomology of a smooth projective variety $Y$ carries a (pure) Hodge structure of weight $k$:
The cohomology of a not necessarily projective and not necessarily smooth varieties carries a mixed Hodge structure (MHS): given such a variety $Y$, this is a pair $(F^{\bullet},W^{\bullet})$ where
When $Y$ is projective, we have $W_{k}H^{k}(Y,\mathbb{Q})=H^{k}(Y,\mathbb{Q})$, and we say that the cohomology of $Y$ is pure if $W_{k-1}(H^{k}(Y,\mathbb{Q}))=0$. A key, defining criterion of purity is the following:
Proposition 4.1.
A projective algebraic variety $Y$ has pure cohomology if and only if for some (and thus for any) birational map $\widetilde{Y}\to Y$ from a smooth projective variety $\widetilde{Y}$ to $Y$, the pullback map $H^{\bullet}(Y)\to H^{\bullet}(\widetilde{Y})$ is injective.
Some extensions of this notion:
  • Compactly supported cohomology $H^{\bullet}_{c}(Y)$ has MHS.
  • MHS may be defined in a natural way on intersection cohomology and compactly supported intersection cohomology as well. Remarkably, when $Y$ is projective, the Hodge structure on $IH^{\bullet}(Y)$ is pure.
  • Finally, one can extend the notion of MHS to intersection cohomology with twisted coefficients (see §1.2), replacing the constant sheaf $\mathbb{Q}_{Y}$ with any local system $\mathcal{L}$ supported on an open dense subset of $Y$.
Then we have the following purity statement (see [25]).
Proposition 4.2.
Let $X$ be a semiprojective variety (see Definition 3.8). Then the intersection cohomology $IH^{\bullet}(X)$ of $X$ has pure Hodge structure.
Corollary 4.3.
The intersection cohomology $IH^{\bullet}(U_{V})$ of the local model $U_{V}$ (see Corollary 3.9) has pure Hodge structure.
4.2. The motivic property
Definition 4.4.
Let $Gr^{W}_{k}(H_{c}^{j}(Y)):=W_{k}(H_{c}^{j}(Y))/W_{k-1}(H_{c}^{j}(Y))\otimes\mathbb{C}$ be the weight-$k$ part of the $j$-th compactly supported cohomology of $Y$. The weight polynomial of $Y$ is the generating function
$E_{t}(Y)=\sum_{j,k}(-1)^{j}\dim Gr^{W}_{k}(H_{c}^{j}(Y))t^{k}.$
Proposition 4.5.
If $Z\subset Y$ is a locally closed subvariety of the variety $Y$, then
$E_{t}(Y)=E_{t}(Z)+E_{t}(Y\setminus Z).$
4.3. MHS and the Decomposition Theorem
Let $f:X\rightarrow Y$ be a proper map of algebraic varieties and recall the notation of Section 1.6. It is a deep statement of Saito’s theory of mixed Hodge modules [9, 6, 33] that it is always possible to choose an isomorphism in the Decomposition Theorem 1.7 which respects the Mixed Hodge structures in (1.7.C) and (1.7.B). In particular, the MHS on the cohomology of the fiber $f^{-1}(p)$ is the direct sum of the MHS on the stalks of intersection cohomology sheaves with twisted coefficients on the right hand side.
In more detail, denote by $\iota:p\hookrightarrow Y$ the inclusion; then base change induces an isomorphism in the derived category of mixed Hodge modules
$\iota^{*}Rf_{*}\mathbb{Q}_{X}\cong R\Gamma(f^{-1}(p),\mathbb{Q}_{X}),$
so that
$H^{k}(f^{-1}(p),\mathbb{Q})\cong H^{k}(p,\iota^{*}Rf_{*}\mathbb{Q}_{X})$
is an isomorphism of MHS. On the other hand, applying the pull-back $\iota^{*}$ to both sides of (1.7.B) and taking cohomology, yields the following direct sum decomposition, respecting the MHS:
(25) $H^{k}(f^{-1}(p),\mathbb{Q})\cong\bigoplus_{\overline{Y}_{\alpha}\ni p}\bigoplus_{d=d_{\alpha}-\boldsymbol{r}(f)}^{d_{\alpha}+\boldsymbol{r}(f)}H^{k-d}\left(p,\iota^{*}IC(\overline{Y}_{\alpha},\mathcal{L}_{\alpha}^{d})\right),$
where $d_{\alpha}=\dim X-\dim Y_{\alpha}$. In particular, if all summands on the right hand side are pure, then the cohomology of the fiber over the point $p$ is pure as well.
However, calculating the induced MHS on the right hand side tends to be very difficult in general, in particular, because the action of the fundamental group of the strata can modify the MHS of the summands in a nontrivial way. Fortunately, in our case, the action of the fundamental group is essentially trivial, and thus we have good control over the relevant MHSs.
We are grateful to Claude Sabbah, who communicated to us the following precise technical statement, supplementing the Decomposition Theorem 1.7.
Proposition 4.6.
Assume the setup of Theorem 1.7.
  1. (1)
    Let $p$ be a point in $Y$ such that for each support stratum $Y_{\alpha}$ containing $p$ in its closure, the following properties holds:
    1. (a)
      (trivial monodromy) There is a neighborhood $D$ of $p$, such that the monodromy of $\mathcal{L}_{\alpha}$ in $D\cap Y_{\alpha}$ is trivial, and hence the local system is constant;
    2. (b)
      for all $d\in\mathbb{Z}$, and all $\alpha$, the Hodge filtration on $\mathcal{L}_{\alpha}^{d}$ in $D\cap Y_{\alpha}$ is trivial of weight $d$;
    3. (c)
      the stalk $\mathcal{H}^{j}\left(IC(Y_{\alpha},\mathbb{Q}_{\overline{Y}_{\alpha}})\right)_{p}$ is pure of weight $j$.
    Then the cohomology of the fiber at $p$ has pure MHS, i.e. $H^{k}(f^{-1}(p),\mathbb{Q})$ is pure of weight $k$.
  2. (2)
    (variant with finite monodromy) The same conclusion holds if (a) is replaced by the following: there exists a finite morphism $g_{\alpha}:\overline{\mathcal{Y}}_{\alpha}\rightarrow\overline{Y}_{\alpha}\cap D$, which induces a finite étale morphism $\mathcal{Y}_{\alpha}:=g_{\alpha}^{-1}(Y_{\alpha}\cap D)\rightarrow Y_{\alpha}\cap D$ such that $g_{\alpha}^{*}\mathcal{L}_{\alpha}^{d}$ is constant on $\mathcal{Y}_{\alpha}$ and the stalk $\mathcal{H}^{j}\left(IC(\overline{\mathcal{Y}}_{\alpha},\mathbb{Q}_{\mathcal{Y}_{\alpha}})\right)_{p^{\prime}}$ is pure of weight $j$ for all $p^{\prime}\in g_{\alpha}^{-1}(p)$.
  3. (3)
    If in part (1) we replace $D\cap Y_{\alpha}$ by $Y_{\alpha}$ in condition (a) above, then (a) implies (b). The analogous statement holds in the finite monodromy case as well (cf. part (2)).
Proof.
(1) This follows from the existence of the pull-back functor $\iota^{*}$ as sketched above. The only thing to observe is that conditions (a) and (b) imply that
$H^{k-d}\left(p,\iota^{*}IC(\overline{Y}_{\alpha},\mathcal{L}_{\alpha}^{d})\right)\cong\mathcal{H}^{k-d}\left(IC(\overline{Y}_{\alpha},\mathcal{L}_{\alpha}^{d})\right)_{p}=\mathcal{H}^{k-d}\left(IC(\overline{Y}_{\alpha},\mathbb{Q}_{{Y}_{\alpha}})\right)_{p}\otimes L^{d}_{\alpha}.$
By our assumptions, the latter is pure of weight $k-d+d=k$.
For part (2), we note that the decomposition theorem for the finite morphism $g_{\alpha}$ implies that $IC(\overline{Y}_{\alpha},\mathcal{L}^{d}_{\alpha})$ is a direct summand of $g_{\alpha,*}IC(\overline{\mathcal{Y}}_{\alpha},g_{\alpha}^{*}\mathcal{L}^{d}_{\alpha})$, hence, for all $p\in\overline{Y}_{\alpha}$ and all $j\in\mathbb{Z}$, $\mathcal{H}^{j}\left(IC(\overline{Y}_{\alpha},\mathcal{L}_{\alpha}^{d})\right)_{p}$ is a direct summand of
$\bigoplus_{p^{\prime}\in g_{\alpha}^{-1}(p)}\mathcal{H}^{j}\left(IC(\overline{\mathcal{Y}}_{\alpha},g_{\alpha}^{*}\mathcal{L}_{\alpha}^{d})\right)_{p^{\prime}}$
as MHS. Since the latter is pure by (1), so is the former.
Finally, for part (3), we note that if $\mathcal{L}^{d}_{\alpha}$ is constant on the quasi-projective variety $Y_{\alpha}$, then the polarizable variation of Hodge structure it underlies is also constant of weight $d$ (see [31, Example 10.1]). Thus $(a)$ implies $(b)$. The argument in the finite monodromy case follows the proof of part (2). ∎
Remark 4.7.
In later parts of the paper, we will verify the conditions of this statement, and thus conclude that the fibers of our map $\pi$ have pure MHS (see Proposition 5.10 and Lemma 6.3). We will, in fact, only need the more ”elementary” part of this proposition, parts (1) and (2), with (3) providing an alternative shortcut in our arguments.